A disk spinning at an initial angular velocity ω decelerates at a constant rate α until it comes to a stop after a time t. If its inital angular velocity is doubled but all else remains the same, how long will it take to stop?

21/2 t

4t

t

2t
Answer #1

initial angular velcity = \omega

deceleration = \alpha

Final angular velocity = 0

0 = \omega - \alpha t

=> t = \frac{\omega }{\alpha }

Case 2

Initial angluar velocity = 2\omega

deceleration = \alpha

Final angular velocity = 0

0 = 2\omega - \alpha time

=>time = \frac {2\omega} {\alpha} = 2t

Time to stop = 2t